# Exploring Tensor Network Algorithms as a Quantum-Inspired Method for Quantum Reservoir Computing

### v1

@@ Abstract
Quantum reservoir computing (QRC) has emerged as a promising hybrid quantum machine learning (QML) method that leverages the complex dynamics of quantum systems and classical machine learning models. Motivated by the development of this new QML method, we explore how quantum-inspired techniques like tensor networks (TNs), specifically the Time Dependent Variational Principle (TDVP) with Matrix Product State (MPS), can be used for the QRC algorithm. To demonstrate the utility of our quantum-inspired method, we performed numerical experiments on the MNIST dataset and compared the performance of our quantum-inspired QRC with different classical machine learning (ML) methods. The results reveal that high-quality embeddings can be generated by performing the time-evolution of MPS system consisting of one-dimensional chain of Rydberg atoms. This quantum-inspired method is highly scalable, enabling the simulation of 100 qubits with a low classical computing overhead. Finally, this study also underscores the potential of tensor networks as quantum-inspired algorithms to enhance the capability of quantum machine learning algorithms to study datasets with large numbers of features.
@@ I Introduction
The convergence of quantum computing and machine learning has ushered in a new era of computational possibilities, collectively known as quantum machine learning (QML) [1, 2, 3, 4, 5]. This interdisciplinary field seeks to exploit the unique capabilities of quantum computers to enhance the performance of many machine learning tasks, potentially addressing problems that are intractable within the context of the classical computing paradigm. Among the various QML approaches, quantum reservoir computing (QRC) [6, 7, 8, 9, 10] has emerged as a particularly promising framework. By harnessing the complex dynamics of quantum systems, QRC offers an innovative method to process information and perform machine learning tasks, thereby circumventing some of the limitations inherent in traditional QML algorithms.
Inspired by classical reservoir computing (RC) [11, 12, 13, 14], QRC leverages the intrinsic dynamics of quantum systems to process data. In classical RC, a fixed dynamical system, known as the reservoir, transforms input data into a high-dimensional space, thus enhancing the features used for training through simple linear regression. QRC extends this concept into the quantum realm, utilizing the exponentially large Hilbert space of quantum systems to generate correlations that are classically intractable.
The promise of many QML algorithms lies in the ability to use near-term quantum devices to produce correlations that would otherwise require exponential classical resources to simulate [1]. However, practical applications of QML are still limited, primarily due to the challenges posed by noisy intermediate-scale quantum devices (NISQ) [15] and the complexity of training quantum models on quantum hardware. Additionally, the availability of quantum computers capable of handling a large number of qubits necessary to encode extensive data features is limited, further hindering the widespread adoption of QML [16, 17]. An alternative approach is to leverage quantum-inspired algorithms based on tensor network (TN) methods [18, 19, 20, 21, 22]. Specifically, tensor network algorithms can simulate the dynamics of certain quantum systems by representing the quantum states and operators in compressed form [23, 24, 25]. This approach significantly reduces the computational resources required, making it feasible to simulate specific quantum systems with a large number of qubits on classical hardware. Due to their utility, TN models have been effectively used as quantum-inspired algorithm to simulate certain quantum circuit models and QML techniques, enabling the exploration of many quantum algorithms using current classical devices [26, 27, 28, 29].
Recent experimental study of the QRC algorithm has shown that this QML method can be implemented effectively on analog quantum hardware up to 108 qubits by encoding classical data to a 1D chain or 2D systems of Rydberg atoms [6]. These results have motivated our exploration of the tensor network algorithm as a quantum-inspired technique for QRC. Specifically, TNs can efficiently simulate quantum state evolutions [23, 30, 31, 32, 33] and compute expectation values [19, 25] for certain quantum systems within a certain approximation of their dynamics [34, 35]. Secondly, tensor networks scale favorably with an increasing number of qubits for systems with low entanglement, mitigating the exponential complexity by exploiting quantum correlations, thus allowing the simulation of large-scale quantum systems with reduced computational overhead [18]. Finally, since the goal of the QRC technique is to generate new features based on the computation of expectation values from certain Hamiltonian dynamics [6], it naturally raises the question if quantum-inspired technique like TNs with MPS can also produce similar ML performance. Such an investigation can provide a practical solution to explore QRC on classical hardware at scale, enabling researchers to investigate QRC properties and performance for datasets with more features, thereby accelerating algorithm development and providing insights into scalability and real-world applicability.
@@
In this paper, we leveraged tensor network algorithms, specifically the time-dependent variational principle (TDVP) [31, 32, 33] and the matrix product state (MPS) to to perform the time-evolution of 1D Rydberg atoms, with the goal of providing heuristic examples of how the QRC technique can be implemented at scale using a quantum-inspired technique. MPS-TDVP techniques based on the Lie-Trotter decomposition have been known to achieve a balance between performance and accuracy when they are used to simulate quantum dynamics such as time evolution [33]. To demonstrate the efficacy of TN, we performed numerical experiments using the MNIST [36] dataset using our MPS-based QRC for classification. Our numerical results suggest that the TDVP and MPS techniques are highly scalable, which opens up more possibilities to apply QRC to datasets with more features using classical computational resources.
@@ II Methodology
II.1 QRC Method with Rydberg Hamiltonian
The QRC algorithm consists of three primary components: data encoding into a Hamiltonian representing certain quantum dynamics, obtaining a quantum embedding through expectation value calculation, and a simple classical machine learning task. The QRC begins by encoding the data features into the parameters of the Rydberg Hamiltonian [38], as described in:
@@ II.1 QRC Method with Rydberg Hamiltonian
In this Hamiltonian, $\Omega_{j}$ represents the Rabi drive amplitude between a ground state (|$\ket{g_{j}}$, with $j$ indexing the atoms) and a highly excited Rydberg state (|$\ket{r_{j}}$). $\Delta_{j}$, the detuning of the driving laser field for atom $j$, and $\phi_{j}$, the laser phase for atom $j$. Furthermore, $V_{jk}$ describes the van der Waals interaction between atoms $j$ and $k$, which can be derived from the geometry of the lattice as $V_{jk}=C/\lVert\mathbf{r}_{j}-\mathbf{r}_{k}\rVert^{6}$. Within our numerical experiments, we used a one-dimensional chain as the lattice geometry with the distance between two atoms being $11\,\mathrm{\mu m}$ and the values of $\Omega=2\pi$ and $\phi=0$. The distance in the range of $10-11\,\mathrm{\mu m}$ yields the best accuracy results for the ML tasks. Other studies in QRC have also suggested distance within this value range for optimal ML performance [6].
@@
Data encoding is achieved through site-dependent local detunings, represented as $\Delta_{j}=x_{j}$ [6]. Consequently, a $N$-qubit system is capable of encoding $N$ features. Following data encoding, the final step involves classical post-processing. To achieve this, we obtain the data embedding vectors, which can be derived from the Hamiltonian dynamics of the quantum system. Specifically, after encoding the data through site-dependent local detunings, represented as $\Delta_{j}=x_{j}$, the $N$-qubit system transitions from an all spin-up ground state under the influence of the specifically designed Rydberg Hamiltonian. The quantum dynamics are then examined over several successive time steps. At each time step, the expectation values of local observables are measured, typically one- and two-point correlators on a computational basis, such as $\langle Z_{j}\rangle$ and $\langle Z_{j}Z_{k}\rangle$. These local observables then form the data embedding vectors, $u_{i}[n]$, with $i$ indexing the different correlators and probe times. These vectors are essential for the classical post-processing step. To obtain these expectation values, the TN algorithm has been used.
@@
II.2 Tensor Network Methods for Dynamical Simulation of Rydberg Hamiltonian for QRC
@@
In the realm of tensor network algorithms, the time-dependent variational principle (TDVP) is a robust method to simulate the dynamics of quantum systems, particularly those with a substantial number of qubits. TDVP can be implemented in two primary forms: single-site TDVP and two-site TDVP. Both approaches are crucial for efficiently evolving quantum states represented by Matrix Product States (MPS) under the influence of a Hamiltonian, represented by the Matrix Product Operator (MPO) format. Single-site TDVP focuses on updating one site (or tensor) of the MPS at a time, ensuring relatively low computational cost and suitability for large systems, while maintaining unitary time evolution, energy conservation [32], and numerical stability [44, 45]. However, single-site TDVP uses a fixed-rank integration scheme, which does not allow the bond dimension to grow during the time evolution, thus limiting its applications to simulate only certain quantum dynamics [35, 46]. In contrast, TDVP for two sites updates pairs of neighboring sites simultaneously, allowing for a more accurate representation of correlations, especially in systems interacting strongly, but has a much higher computational cost [31].
@@
For the simulation of the quantum system in this study, the TeNPy library [39] was utilized. TeNPy is a versatile and efficient library for implementing tensor network algorithms, including single-site and two-site TDVP. It provides robust tools for handling MPS and MPO representations, making it ideal for simulating quantum system evolution. In this simulation, we encoded the classical data in a 1D chain. The Rydberg Hamiltonian was represented using the MPO format, and the quantum system, initially set in an all-up state,was evolved using the TDVP algorithm. Correlators were computed at each time step to generate the nonlinear embedding.
@@
It is important to note that the interaction term $V_{jk}$ decreases very rapidly as the distance between the atoms increases. Consequently, $V_{jk}$ becomes negligible for the farthest neighbors, which can be safely ignored in the calculations. This rapid decrease in interaction strength allows us to reduce the dimensions of MPOs, leading to a significant reduction in computational time. Furthermore, we can optimize data embedding by excluding expectation values $\langle Z_{j}Z_{k}\rangle$ for pairs of sites $j$ and $k$ that do not interact in the Hamiltonian. By focusing only on the relevant interactions, we can streamline our data processing and further enhance computational efficiency. For instance, if we consider the case of 25 qubits, the original number of correlators for a particular time step would be 325. However, by choosing a truncation limit for $V_{ij}$ of $10^{-4}$, the number of correlators can be reduced to 247.
@@ II.3 Data Preprocessing, QRC Simulation, and Classical Machine Learning Methods
To investigate the efficacy of our quantum-inspired QRC technique, we applied the method to perform a classification task using the MNIST dataset [36]. The initial step in our methodology involves preprocessing the MNIST dataset, which consists of 60,000 training samples and 10,000 test samples, each represented as grayscale images of $28\times 28$ pixels. Principal Component Analysis (PCA) [40] was used to reduce the dimensionality of the data to facilitate the embedding process. Each component of the PCA-reduced data was mapped to a qubit as described in the methodology.

### v3

# Harnessing Quantum Dynamics for Robust and Scalable Quantum Extreme Learning Machines

@@ Abstract
Quantum Extreme Learning Machine (QELM) is an emerging hybrid quantum machine learning framework that leverages quantum system dynamics to enhance classical models. However, QELM can suffer from the exponential concentration problem, where excessive entanglement reduces model expressivity. In this work, we gain insight into this challenge and demonstrate how tensor network methods specifically, the Time Dependent Variational Principle (TDVP) with Matrix Product States (MPS) can efficiently simulate quantum systems while controlling entanglement and mitigating exponential concentration. Using numerical experiments on the Modified National Institute of Standards and Technology (MNIST) dataset, we show that time-evolving an MPS system modeled as a chain of Rydberg atoms produces high-quality data embeddings with low classical computational overhead. Our findings indicate that exact simulation of quantum dynamics is not necessary for strong machine learning performance; even approximate quantum embeddings can yield competitive results. Furthermore, we observe that both increased disorder in the quantum state achieved by tuning Hamiltonian parameters and careful control of entanglement directly correlate with improved model accuracy, highlighting the importance of these factors in optimizing QELM performance.
@@ I Introduction
The intersection of quantum computing and machine learning (ML) has catalyzed the emergence of quantum machine learning (QML), a rapidly evolving field that seeks to harness quantum phenomena to enhance classical data processing and pattern recognition tasks [1, 2, 3, 4, 5]. Among the varied approaches within QML, hybrid frameworks that integrate quantum system dynamics with classical models such as Quantum Reservoir Computing (QRC) and Quantum Extreme Learning Machines (QELM) [6, 7, 8, 9, 10, 11, 12] have shown particular promise. These methods leverage the complex, high-dimensional dynamics of quantum systems to transform input data into expressive feature spaces, potentially enabling efficient solutions to problems that are intractable for classical algorithms.
Inspired by classical reservoir computing (RC) and Extreme Learning Machines (ELM)[13, 14, 15, 16, 17, 18, 19], QRC and QELM leverage the intrinsic dynamics of quantum systems to process data. In classical RC and ELM, a fixed dynamical system, known as the reservoir, transforms input data into a high-dimensional space, thus enhancing the features used for training through simple linear regression. Although RC and ELM seemingly similar, a main difference resides in the fact that RC exploits the natural dynamics of the substrate (reservoir) as an internal memory of past input information while ELM does not [10]. This is why RC is reserved for time series prediction, whereas ELM is used for tasks such as classification and regression. QRC and QELM extend this concept into the quantum realm, utilizing the exponentially large Hilbert space of quantum systems to generate correlations that are classically intractable. Recent experimental study of the QRC algorithm has shown that this QML method can be implemented effectively on analog quantum hardware up to 108 qubits by encoding classical data into different geometrical structures of Rydberg atoms [6]. Beyond Rydberg Hamiltonians, many studies have investigated the interplay between the quantum encoding layer and the Hamiltonian used to generate the quantum reservoir, and how this interaction can influence subsequent ML tasks. [20, 21, 22, 23, 24, 25].
Although QELM shows great potential, recent literature points out major challenges, notably the risk of exponential concentration. This issue arises when expectation value measurements cluster around specific values regardless of the input, resulting in nearly indistinguishable feature representations and ultimately impairing the effectiveness of downstream classical ML algorithms. Such a phenomenon is intrinsically linked to excessive entanglement, originated from the quantum feature encoding scheme or the Hamiltonian evolution generating the quantum reservoir [11]. As a result, understanding the physics to carefully manage the entanglement growth within the quantum dynamics is essential for optimal performance in QELM.
Among many quantum many-body systems, the 1D Rydberg model exhibits rich quantum phases by tuning internal parameters within the Hamiltonian such as the Rabi frequency, detuning, as well the distance between the atoms within the chain [26, 27, 28, 29]. In addition, QRC and QELM have been shown to be viable on system like Rydberg based quantum computer [6, 30, 31]. Within this study, we leverage the quantum physics of the 1D Rydberg chain to optimize the performance of our QELM model by tuning the different parameters within the Hamiltonian to produce optimal quantum features.
Beyond exact simulation of quantum dynamics, quantum-inspired algorithms based on tensor network (TN) methods have emerged as an important tool to study quantum many-body systems including Rydberg-based models [32, 33, 34, 35, 36, 37, 38, 39]. TN algorithms offer a key advantage: they efficiently represent and manipulate quantum states with limited entanglement, enabling the simulation of large quantum systems on classical hardware by compressing the state space [40, 41, 42]. Specifically, TNs can efficiently simulate quantum state evolutions [40, 43, 44, 45, 46, 47] and compute expectation values [33, 42] for certain quantum systems within a certain approximation of their dynamics [48, 49]. Secondly, TNs scale favorably with an increasing number of qubits for systems with low entanglement, mitigating the exponential complexity by exploiting quantum correlations, thus allowing the simulation of large-scale quantum systems with reduced computational overhead [32]. Finally, since our quantum reservoir is generated by evolution over time, the nonequilibrium state after the evolution can result in entanglement spreading throughout the system [50, 51], resulting in lower performance of the ML model due to the problem of exponential concentration. Consequently, compression using TNs not only reduces computational resources, but also provides a means of controlling the entanglement structure which is particularly relevant for addressing the exponential concentration problem observed in QELM [11]. Thus, TN methods present a practical and scalable pathway to advance QELM and other quantum-inspired machine learning frameworks on classical devices.
@@
In this paper, we leveraged TN algorithms, specifically the time-dependent variational principle (TDVP) [44, 45, 46] and the matrix product state (MPS) to perform the time-evolution of 1D Rydberg atoms, with the goal of providing an example of how the QELM technique can be optimized and implemented at scale using a quantum-inspired technique. MPS-TDVP techniques based on the Lie-Trotter decomposition have been known to achieve a balance between performance and accuracy when they are used to simulate quantum dynamics such as time evolution [46]. Specifically, TDVP restricts dynamics of the many-body wavefunction within a variational MPS manifold during the time evolution [49]. As a result, this can provide a pathway to manage the entanglement growth during the time evolution. To demonstrate the efficacy of TNs, we performed numerical experiments using the MNIST [52] dataset using our MPS-based QELM for classification. Our results suggest that the TDVP and MPS techniques are highly scalable, which opens up more possibilities to apply QELM to datasets with more features using classical computational resources.
In short, our experiments show that the TDVP-MPS approach enables efficient simulation of quantum dynamics for systems with a higher number of qubits on classical hardware, with computational overhead that scales favorably. We find that achieving strong ML performance does not require highly precise simulation of quantum dynamics; even approximate embeddings generated by scalable TN algorithms yield classification accuracies comparable to those of nonlinear classical models, including neural networks (NNs). Furthermore, the degree of disorder in the quantum state controlled by tuning Hamiltonian parameters directly correlates with improved model performance. Increased disorder enhances the diversity and expressiveness of the quantum embeddings, leading to higher classification accuracy. Analysis of quantum entanglement reveals that, while some entanglement is necessary to enrich the feature space, excessive entanglement can result in feature concentration and diminished model performance.
Figure 1: A schematic representation of our tensor network approach for QELM with Rydberg Hamiltoninan embedding and how the different parameters in the Hamiltonian can be used to control entanglement and disorder in the Rydberg dynamics. (a) The workflow includes data preprocessing, constructing the embedding by computing the expectation values of correlators with tensor network formalism at each time step dtdt up to total time $T$, and finally performing prediction using classical machine learning regression methods. (b) The figure also demonstrates the optimization of Hamiltonian parameters, such as the Rabi frequency ($\Omega$) and the distance between two atoms. The highlighted regions represent parameters which yield optimal ML performance. By fixing one parameter and varying the other, the system’s disorder can be maximized, which enhances machine learning accuracy as greater disorder generally leads to better performance. Additionally, the effect of entanglement entropy is shown, indicating that a moderate amount of entanglement is optimal for the algorithm; both too little and too much entanglement can be detrimental. This data was generated using the TDVP two-site algorithm.
@@ II Methodology
A schematic workflow of our TN simulation of Rydberg Hamiltonian dynamics to generate the new embedding for ML task is shown in Fig.(1a). In general, a QELM algorithm consists of three primary components: data encoding using quantum feature map, performing Hamiltonian simulation with time evolution to increase the feature space; obtaining a quantum embedding through expectation value calculation; and training a simple classical ML model using the new quantum embedding. In our approach, we utilized a Rydberg Hamiltonian and its dynamics to encode the data as well as to generate the new quantum embeddings. In addition, rather than performing exact simulation of the quantum dynamics, TN methods like MPS and TDVP methods were used to perform the encoding and time-evolution in a 1D Rydberg chain.
II.1 Rydberg Hamiltonian Encoding and Quantum Embedding generation
This method begins by encoding the data features into the parameters of the Rydberg Hamiltonian [53], as described in:
@@ II.1 Rydberg Hamiltonian Encoding and Quantum Embedding generation
[…]
@@
Data encoding is achieved through site-dependent local detunings, represented as $\Delta_{j}=x_{j}$ [6]. Consequently, a $N$-qubit system is capable of encoding $N$ features. Once the detuning values are fixed through data encoding, the remaining Hamiltonian parameters to set are the Rabi frequency $\Omega$ and the interatomic distance $d$, which determines the interaction strength $V_{jk}$.The selection and influence of these parameters are discussed in detail in the Results and Discussion section. As illustrated in Fig.(1b), optimal model performance is achieved by tuning the Hamiltonian parameters to maximize disorder in the quantum states while maintaining moderate entanglement. Following data encoding, the final step involves classical post-processing. To achieve this, we obtain the data embedding vectors, which can be derived from the Hamiltonian dynamics of the quantum system. Specifically, after encoding the data through site-dependent local detunings, represented as $\Delta_{j}=x_{j}$, the $N$-qubit system transitions from an all spin-up ground state under the influence of the specifically designed Rydberg Hamiltonian. The quantum dynamics are then examined over several successive time steps. At each time step, the expectation values of local observables are measured, typically one- and two-point correlators on a computational basis, such as $\langle Z_{j}\rangle$ and $\langle Z_{j}Z_{k}\rangle$. These local observables then form the data embedding vectors, $u_{i}[n]$, with $i$ indexing the different correlators and probe times. These vectors are essential for the classical post-processing step. To obtain these expectation values, the TN algorithm has been used.
@@
II.2 Tensor Network Methods for Dynamical Simulation of Rydberg Hamiltonian for QELM
@@
In the realm of TN algorithms, the TDVP is a robust method to simulate the dynamics of quantum systems, particularly those with a substantial number of qubits. TDVP can be implemented in two primary forms: single-site TDVP and two-site TDVP. Both approaches are crucial for efficiently evolving quantum states represented by MPS under the influence of a Hamiltonian, represented by the Matrix Product Operator (MPO) format. Single-site TDVP focuses on updating one site (or tensor) of the MPS at a time, achieving relatively low computational cost and suitability for large systems while maintaining unitary time evolution, energy conservation [45], and numerical stability [54, 55]. However, single-site TDVP uses a fixed-rank integration scheme, which does not allow the bond dimension to grow during the time evolution, limiting its applications to simulate only certain quantum dynamics [49, 56]. In contrast, TDVP for two sites updates pairs of neighboring sites simultaneously, allowing for a more accurate representation of correlations, especially in strongly interacting systems, but has a much higher computational cost [44].
@@
For the simulation of the quantum system in this study, the TeNPy library [57] was utilized. TeNPy is a versatile and efficient library for implementing TN algorithms, including single-site and two-site TDVP. It provides robust tools for handling MPS and MPO representations, making it ideal for simulating quantum system evolution. In this simulation, we encoded the classical data in a 1D chain. The Rydberg Hamiltonian was represented using the MPO format, and the quantum system, initially set in an all-up state,was evolved using the TDVP algorithm. Correlators were computed at each time step to generate the nonlinear embedding.
@@
It is important to note that the interaction term $V_{jk}$ decreases very rapidly as the distance between the atoms increases. Consequently, $V_{jk}$ becomes negligible for the farthest neighbors, which can be safely ignored in the calculations. This rapid decrease in interaction strength allows us to reduce the dimensions of MPOs, leading to a significant reduction in computational time. Furthermore, we can optimize data embedding by excluding expectation values $\langle Z_{j}Z_{k}\rangle$ for pairs of sites $j$ and $k$ that do not interact in the Hamiltonian. By focusing only on the relevant interactions, we can streamline our data processing and further enhance computational efficiency. For instance, if we consider the case of 25 qubits, the original number of correlators for a particular time step would be 325. However, by choosing a truncation limit for $V_{jk}$ of $10^{-4}$, the number of correlators can be reduced to 247. The detailed analysis of $V_{jk}$ threshold is given in [58].
@@ II.3 Data Preprocessing, QRC Simulation, and Classical Machine Learning Methods
To investigate the efficacy of our quantum-inspired QRC technique, we applied the method to perform a classification task using the MNIST dataset [52]. The initial step in our methodology involves preprocessing the MNIST dataset, which consists of 60,000 training samples and 10,000 test samples, each represented as grayscale images of $28\times 28$ pixels. Principal Component Analysis (PCA) [59] was used to reduce the dimensionality of the data to facilitate the embedding process. Additionally, we rescaled the $x_{j}$ inputs to the range [6,6][-6,6]. Each component of the PCA-reduced data was mapped to a qubit as described in the methodology.
Figure 2: This plot illustrates the variation of entanglement entropy with increasing time for different values of $\Omega$. This plot was generated using TDVP two site algorithm.